# Reconciling aging and slip state evolutions from laboratory-derived canons of rate-and-state frictionF

### v2

@@ 5 Discussion
As shown in §4, the slip law is the very requirement from VS and SS tests, while the aging law is one of the natural functional forms satisfying the requirements from SHS and SS tests. However, neither law can simultaneously satisfy requirements from VS, SHS (SHS||||NSC, where behaviors contradict the VS requirement), and SS tests. In fact, no differential equation can simultaneously reproduce the strengthening rate observed in VS tests and that observed in SHS||||NSC tests (§4.1). This simple mathematical fact is, we would say, the fundamental reason why we have been struggling for so long to reconcile the slip and aging laws (e.g., Kato & Tullis, 2001). This issue has critical geophysical implications because the two laws predict quite different behaviors for many aspects of earthquake cycles, including the recurrence interval (Kato & Hirasawa, 1999), the earthquake nucleation (Ampuero & Rubin, 2008), and the aseismic interaction between megathrust earthquakes and repeaters (Ariyoshi et al., 2014).
The present work has theoretically shown that simultaneous reproduction of the VS and SHS canons necessarily requires switching between the aging-law-like evolution and the slip-law-like evolution, while evolution laws involving such a switch are to predict behaviors violating the VS canon and the SHS canon under the conditions where the crossover occurs. Accepting such non-desired behaviors predicted as inevitable side effects, we carried on our development because behaviors violating the VS and SHS canons near the Slip/Aging switch are expected to occur in the $(V,\theta)$ range not well probed experimentally so far and hence can be hidden there. However, given this acceptance of the yet-to-be-observed side effects, we still must be aware that only certain forms of aging-slip switches are consistent with the existing experimental results. For example, the aging-slip switch introduced in the composite law (Kato & Tullis, 2001) violates the SS canon in the range already confirmed by experiments, not only having the side effects (§4.1) violating the VS and SHS canons at the crossover regime.
@@
In this §5, we work on several issues arising from the above-summarized insights obtained by §4. The first issue is whether the control factor of the canons-required aging-slip switch must be $\Omega$. In §5.1, we further generalize the derivations of §4.2.2 to find that the $\Omega$-dependent switch is nearly a universal requirement by the three canons. The second issue is the physical meaning of the $\Omega$-dependent switch of the evolution mechanism. In §5.2, we discuss this point from a macrophysical viewpoint. The third issue is experimental validation. Although the presently proposed class of evolution laws involving the $\Omega$-dependent switch, deduced from the compilation of the three canons without heuristics, is nearly the unique solution, this does not guarantee the predicted behaviors, especially the side effects of the switching, are the case. In §5.3, we will validate the modified composite law (eq. 96) using an existing experiment that seems to include the $(V,\theta)$ range where the predicted side effects of the aging-slip switch begin to appear.
@@
However, even the generalized version (eq. 86) was derived under two a priori assumptions besides the canons (§4.2.2). The first assumption was the TDF, as mentioned already. It represents an a priori belief (as a working hypothesis) that variation in strength consists of a response to the time-lapse ($F_{t}$) and that to the slip ($F_{\delta}$). In the derivation, the TDF contributed substantially to assigning the VS and SHS canons to the slip response ($f_{\delta}$) and the time response ($f_{t}$), respectively. By contrast, the derivation below shall be done without the TDF assumption. The second assumption was the functional limitation $\dot{\theta}=f(V,\theta)$. We will keep this. One reason is to forbid using an almighty oracle variable for switching evolution functions between VS and SHS||||NSC tests; the use of an oracle would be mere tautology of experimental constraints shown in §3. Although oracle exclusion may be achievable in other ways, we use $\dot{\theta}=f(V,\theta)$, which has been standard since Ruina (1983), for simplicity of derivation. The functional space $\dot{\theta}=f(V,\theta)$ contains $\dot{\theta}=f(V,\theta,\tau)$ as $\tau$ can be expressed in terms of $V$ and $\theta$ (eqs. 1 or 17). Still, we have to admit that the $f(V,\theta)$ presumption has precluded some plausible effects, such as the shear stress dependence (Nagata et al., 2012), written as $\dot{\theta}=f(V,\theta,\dot{\tau})$ [or equivalently, $\dot{\theta}=f(V,\theta,\dot{V})$]. Considerations of evolution functions not being of $f(V,\theta)$ are left to future studies.
Now, we start the derivation of a further generalized form of eq. (86) with the TDF assumption removed, whereas we still assume $\dot{\theta}=f(V,\theta)$ as mentioned above. We only consider $\theta,\theta_{\rm SS}(V)\gg\theta_{\rm X}$ because the strength minimum can be incorporated by the procedure shown in §4.5. We redo the derivation of experimental constraints from the canonical behaviors without assuming the TDF this time. Constraints from SS requirements remain eq. (31) because the original derivation of eq. (31) in §3.1 did not utilize the TDF.
@@
To express constraints from SHS tests under stationary contact, we introduced the NSC in §3.3. However, the definition of the NSC (eq. 46) already assumes the TDF; hence, we here need to express the stationary contact in other ways. We can do this by replacing ‘NSC’ with the literal definition of stationary contact, ‘$\delta\to 0$’. Repeating §3.3.2 with ‘NSC’ replaced by ‘$\delta\to 0$’, we have
@@
Next, we need to specify the regimes where each constraint applies. Again, we cannot use expressions that can be defined only under the TDF. As discussed in §4.2, VS tests require (i) $f=f_{\rm VS}$ for $\Omega\gtrsim V_{\rm after}/V_{\rm before}$ probed by existing VS tests. Note that this requirement (i) is not conditioned on $V_{\rm after}$ itself because the VS canon has been confirmed over a wide range of absolute slip rates. Because of (i), the canonical behavior of SHS tests ($f=1$, eq. 124) is allowed only in the $V$-$\theta$ regime sufficiently below the steady state that excludes the $\Omega$ range in the strengthening phase following the negative velocity step. Meanwhile, the SHS canon requires (ii) $f=f_{\rm SHS|\delta\to 0}$ for stationary contact. This time, as mentioned earlier, we cannot express stationary contact as the NSC (eq. 46), which is only definable under the TDF. Instead, we here express stationary contact by invoking constant-$\tau$ SHS tests. If we, upon the start of the hold period, drop the shear stress by much more than $A$ from the level [$\tau_{\rm SS}(V_{\rm prior})$] just before the start of hold, stationary contact is realized and kept for a long time throughout the hold period (§A.1). When we consider a large-stress-drop limit of $e^{-\Delta\tau/A}\to 0$, $\Omega\to 0$ holds for $0<t_{\rm h}<\infty$. Therefore, from (ii), we can state $f=f_{\rm SHS|\delta\to 0}$ when $\Omega\to 0$. Thus, from (i) and (ii), we have
@@
The same discussion as that for eq. (86) (+ eq. 72) in §4.2.2 guarantees evolution laws conforming eq. (125) automatically satisfies the SS constraint (eq. 31). Hence, like eq. (86) (+ eq. 72) under the TDF, but without the TDF assumption this time, eq. (125) is a necessary and sufficient condition to reproduce the three canons simultaneously (except for the range not experimentally probed so far).
@@
Below is a cautious remark regarding the difference between the low-velocity limit $V\to 0$ and the far-below-the-steady-state limit $\Omega\to 0$. Because $V\to 0$ and $\Omega\to 0$ are equivalent unless $\theta\to\infty$, we cannot distinguish $V\to 0$ from $\Omega\to 0$ when we consider constant-$\tau$ SHS tests for finite hold time. Therefore, the SHS condition (ii) used in the above derivation of eq. (125), which we took as a constraint at the low-$\Omega$ limit, might be a condition at the low-$V$ limit in fact. This difference may matter in describing the state evolution in loadpoint-hold SHS tests where both $\theta$ and $V$ change gradually, and $\Omega$ is nearly constant so that the limits represented by $\Omega\to 0$ and $V\to 0$ can be different. Hence, some ambiguity is left in the above derivation and eq. (125). Such ambiguity did not exist in the derivation of eq. (86) in §4.2.2 because the NSC cannot exist outside the low-$\Omega$ regime (proven at the beginning of §4). This ambiguity is part of the reason why we first showed eq. (86). Yet, experiments so far confirm that the VS canon holds independently of $V$, which singly rejects the emergence of the asymptotic limits according to the value of $V$, supporting the $\Omega$-dependent switch as expressed in eq. (125). Equation (86) expresses this with a single constant $\beta$. Such $\beta$, as the scale of the aging-slip switch (cf. Appendix F), also exists for eq (125), which is not limited to the TDF (though $\dot{\theta}=f(V,\theta)$ is still assumed), although $f$ in eq (125) is largely unconstrained over a transitional $\Omega$ range.
@@
To sum up, the aging-slip crossover according to $\Omega$ (eq. 125), or the existence of a characteristic value of $\Omega$, $\beta$, in eq. (86), is a general conclusion deduced from the three canons as far as $\dot{\theta}=f(V,\theta)$. Excepting Bhattacharya et al. (2015; 2022) discussed later in §5.3, nobody has observed this crossover, but there is no evidence denying it so far. Given the generality kept in the derivations, if future experiments reject the existence of $\beta$, it will logically mean that the three canons are not the case somewhere in the regimes experimentally probed so far (again, if we ignore the possibility $\dot{\theta}\neq f(\theta,V)$). Simultaneous reproduction of the three canons is the target of the present work and the historical development of the RSF. As has turned out in the present work, the difficulty of reconciling the aging and slip laws is not such a superficial matter that just nobody has yet thought of a clever formula to reproduce the three canons simultaneously. Thence, we believe the general conclusion of this work, that is, aging-slip crossover around $\Omega\sim\beta$, has fundamental importance, warranting future experimental and microphysical studies to verify or falsify it.
@@ 5.2 Macrophysical implications of the $\Omega$-controlled aging-slip switch
Noting $\Omega=\theta/\theta_{\rm SS}$, the $\Omega$-dependent nature of the aging-slip switch means that the switch occurs according to the relative difference of instantaneous $\theta$ from its steady-state value for the instantaneous slip rate. When the strength minimum is incorporated, the switch occurs according to $\tilde{\Omega}/\tilde{\Omega}_{\rm SS}=\tilde{\theta}/\tilde{\theta}_{\rm SS}$, as in $\tilde{\psi}_{\rm c}(\tilde{\Omega}/\tilde{\Omega}_{\rm SS})$ (eq. 116), so that the control factor of the switch becomes the relative difference of $\tilde{\theta}=\theta+\theta_{\rm X}$ from its steady-state value. Then, under Dieterich (1978)’s interpretation of $\theta$ as the average contact time, it is implied the aging-slip switch is controlled by the contact time for $\theta\gg\theta_{\rm X}$, while it becomes almost independent of the contact time for $\theta\ll\theta_{\rm X}$. Such clumsiness, however, disappears in the $\Phi$-notation. From eqs. (15), (105), and (110), we have
@@
For $\theta,\theta_{\rm SS}\gg\theta_{\rm X}$, from eqs. (6) and (31), we have
@@
Hence, for any values of $\theta$, our evolution laws set the aging-slip switch according to $\Phi-\Phi_{\rm SS}$. The threshold $\Phi-\Phi_{\rm SS}$ value of the aging-slip switch is $B\ln\beta$ (eqs. 86 and 126). Note that eq. (126) also indicates that $\tilde{\Omega}/\tilde{\Omega}_{\rm SS}$ affects strength change only through its logarithm. Then, the crossover regime of the aging-slip switch around $\Omega\sim\beta$ in eq. (86) is narrowed in the $\Phi$ scale so that eq. (86) switches the constraints on the evolution depending on whether the strength $\Phi$ is significantly higher or lower than the threshold value $\Phi_{\rm SS}+B\ln\beta$.
Furthermore, noticing that the direct effect does not contribute to the difference of the shear stress from its steady-state value (eq. 7), we see $\tau-\tau_{\rm SS}=\Phi-\Phi_{\rm SS}$. Then, eq. (126) can be rewritten into
@@
Moreover, for interfaces with $|A-B|/B\ll 1$ as typically observed in experiments [e.g., $|A-B|/B\simeq 1/30$ reported by Bhattacharya et al. (2015)], we can propose a more physically intelligible interpretation of the $\Omega$-dependent aging-slip switch as shown below. In the remaining part of this subsection, we assume $|A-B|/B\ll 1$ unless otherwise noted. Then, the steady-state frictional resistance $\tau_{\rm SS}$ remains close to the reference value $\tau_{*}$ unless hugely extensive range of velocities is considered; we obtain the following relation from the SS canon (eq. 26), using the identity eq. (128):
@@
As mentioned in §4.1, Kato & Tullis (2001) proposed to switch the aging and slip laws according to the slip rate $V$, under the expectation that there must be a critical slip rate $V_{\rm c}$ where the crossover occurs. Upon introducing the composite law, Kato & Tullis (2001) said, “To accommodate a friction law to experimental observations of both slide-hold-slide tests and velocity stepping tests, the evolution equation of $\theta$ must be similar to the slip law near the steady state and must approach unity at very small velocities.” [Note: for readability in the present discussion, we have replaced some expressions from the original: state variable $\to\theta$; (3) $\to$ the slip law]. We can then summarize their requirement on the evolution law, here denoted by $f_{\rm KT}$, in a form analogous to eq. (125):
@@
Here, $f_{\rm VS}$ intends, as in eq. (125), the exponential convergence toward the steady state over a constant slip distance (e.g., eq. 43), but does not include the enforcement of steady-state value following the SS canon. We think this is what Kato & Tullis (2001) intended when they mentioned ‘similar to the slip law’ in the above excerpt. Although ‘the slip law’ (eq. 5) enforces both behaviors, it is clear that the latter (SS-canon enforcement) was not intended by them, given that their proposed composite law does not follow the SS canon.
Because $V\to 0$ and $\Omega\to 0$ cannot be distinguished for finite $\theta$ in our derivation of eq. (125), we may say that Kato and Tullis’s (2001) idea expressed by eq. (130) is correct from the conceptual viewpoint of the aging-slip switch. The only major failure of the composite law (eq. 74) is the deviation from the SS canon. [Strictly speaking, eq. (130) has one more problem; constraint by $f_{\rm VS}$ needs to be imposed for a broader $\Omega$ range than said in the first line of eq. (130). This problem can be understood if one notices that the aging law, which does not reproduce the VS canon, actually satisfies eq. (130). Nevertheless, we here ignore this shortcoming for this discussion at an upper conceptual level.]
@@
Leaving aside the workings of the $\Omega$-dependent switch, how can we interpret the physical origin of the $\Omega$-dependent switch? We showed that the $\Omega$-dependent switch can be approximated as a $\tau$-dependent switch for $|A/B-1|\ll 1$. This fact implies a possibility that the aging/slip switch is inherently, rather than apparently, $\tau$-dependent, while the three canons hold only approximately instead. Note, however, that the $\tau$-dependent switch interpretation brings bumps (resembling bumps in Kato and Tullis’s composite law) in SS tests, though only at extremely low $\tau$.
Here, we focus on the interpretation of the physical cause of the $\Omega$-dependent switch itself. Possible microphysics would be countless at this stage. Instead, we point out an inevitability from macrophysics. One feature of the slip law is the symmetric evolution following positive and negative velocity steps, with the same $D_{\rm c}$ (Ampuero & Rubin, 2008). Given that fracture energy is $d\Phi_{\rm SS}(=Bd\ln V)$ multiplied by $D_{\rm c}$, the slip law means the fracture energy is the same for the forward path and the reverse path in the VS canon. Hence, the fracture energy along a closed thermodynamical path in a canonical VS test, a velocity-step series finally returning to its initial slip rate with sufficiently high stiffness, is always zero, qualifying as ‘surface energy’ (Griffith, 1921), which is the fracture energy under quasi-equilibrium (more precisely, quasi-steady states). Besides, thermodynamics tells us that the fracture energy for a far-from-equilibrium path is necessarily more than the quasi-equilibrium fracture energy (Rice, 1978). Therefore, the absolute value of the fracture energy in a large negative VS, experiencing the non-equilibrium trajectory far below the steady state, should be less than the absolute value of the surface energy, while the fracture energy in a large positive VS should be greater than the surface energy. This thermodynamical requirement can be satisfied by the aging effect, which increases the fracture energy (decreases its absolute value) in a large negative VS. Recent observations that $D_{\rm c}$ in the resliding phase of SHS tests increase with $t_{\rm h}$ (Beeler et al., 2022) may be another manifestation of the thermodynamical expectation mentioned above because the resliding phase of SHS can be regarded as a large positive velocity step. Under the above insight from irreversible thermodynamics, $\beta$ can be interpreted to represent the border between the quasi-equilibrium and non-equilibrium (strictly speaking, quasi-steady-state and far-from-steady-state) regimes. The deviation from the slip law is a strict order by thermodynamics, and the border $\beta$ bounding the aging regime will be one example.

### v3

@@ 5 Discussion
As shown in §3, the slip law is the very requirement from VS and SS tests, while the aging law is one of the natural functional forms satisfying the requirements from SHS and SS tests. However, neither law can simultaneously satisfy requirements from VS, SHS (SHS||||NSC, where behaviors contradict the VS canon), and SS tests. In fact, no differential equation can simultaneously reproduce the strengthening rate observed in VS tests and that observed in SHS||||NSC tests (§4.1). This simple mathematical fact is, we would say, the fundamental reason why we have been struggling for so long to reconcile the slip and aging laws (e.g., Kato & Tullis, 2001). This issue has critical geophysical implications because the two laws predict quite different behaviors for many aspects of earthquake cycles that may entail both slip weakening during fault rupture and time-dependent healing of locked zones. Examples include the recurrence interval (Kato & Hirasawa, 1999), the earthquake nucleation (Ampuero & Rubin, 2008), and the aseismic interaction between megathrust earthquakes and repeaters (Ariyoshi et al., 2014).
The present work has theoretically shown that simultaneous reproduction of the VS and SHS canons necessarily requires switching between the aging-law-like evolution and the slip-law-like evolution, while evolution laws involving such a switch are to predict behaviors violating the VS canon and the SHS canon under the conditions where the switch happens. Accepting such non-desired behaviors predicted as inevitable side effects, we carried on our development because behaviors violating the VS and SHS canons near the Slip/Aging switch are expected to occur in the $(V,\theta)$ range not well probed experimentally so far and hence can be hidden there. However, given this acceptance of the yet-to-be-observed side effects, we still must be aware that only certain forms of aging-slip switches are consistent with the existing experimental results. For example, the aging-slip switch introduced in the composite law (Kato & Tullis, 2001) violates the SS canon in the range already confirmed by experiments, not only having the side effects (§4.1) violating the VS and SHS canons at the transient regime.
@@
In this §5, we work on several issues arising from the above-summarized insights obtained by §4. The first issue is whether the control factor of the canons-required aging-slip switch must be $\Omega$. In §5.1, we further generalize the derivations of §4.2.2 to find that the $\Omega$-dependent switch is nearly an inevitable consequence of the three canons. The second issue is the physical meaning of the $\Omega$-dependent switch of the evolution mechanism. In §5.2, we discuss this point from a macrophysical viewpoint. The third issue is experimental validation. Although the presently proposed class of evolution laws involving the $\Omega$-dependent switch, deduced from the compilation of the three canons without heuristics, is nearly the unique solution, this does not guarantee the predicted behaviors, especially the side effects of the switching, are the case. In §5.3, we will validate the modified composite law (eq. 96) using an existing experiment that seems to include the $(V,\theta)$ range where the predicted side effects of the aging-slip switch begin to appear.
@@
However, even the generalized version (eq. 86) was derived under two a priori assumptions besides the canons (§4.2.2). The first assumption was the TDF, as mentioned already. It represents an a priori belief (as a working hypothesis) that variation in strength consists of a response to the time-lapse ($F_{t}$) and that to the slip ($F_{\delta}$). In the derivation, the TDF contributed substantially to assigning the VS and SHS canons to the slip response ($f_{\delta}$) and the time response ($f_{t}$), respectively. By contrast, the derivation below shall be done without the TDF assumption. The second assumption was the functional limitation $\dot{\theta}=f(V,\theta)$. We will keep this. One reason is to forbid using an almighty oracle variable for switching evolution functions between VS and SHS||||NSC tests; the use of an oracle would be mere tautology of experimental constraints shown in §3. Although oracle exclusion may be achievable in other ways, we use $\dot{\theta}=f(V,\theta)$, which has been standard since Ruina (1983), for simplicity of derivation. The functional space $\dot{\theta}=f(V,\theta)$ contains $\dot{\theta}=f(V,\theta,\tau)$ as $\tau$ can be expressed in terms of $V$ and $\theta$ (eq. 1 or 17). Still, we have to admit that the $f(V,\theta)$ presumption has precluded some plausible effects, such as the shear stress dependence (Nagata et al., 2012), written as $\dot{\theta}=f(V,\theta,\dot{\tau})$ [or equivalently, $\dot{\theta}=f(V,\theta,\dot{V})$]. Considerations of evolution functions not being of $f(V,\theta)$ are left to future studies.
Now, we start the derivation of a further generalized form of eq. (86) with the TDF assumption removed, whereas we still assume $\dot{\theta}=f(V,\theta)$ as mentioned above. We only consider $\theta,\theta_{\rm SS}(V)\gg\theta_{\rm X}$ because the strength minimum can be incorporated by the procedure shown in §4.5. We redo the derivation of experimental constraints from the canonical behaviors without assuming the TDF this time. Constraints from the SS canon remain eq. (31) because the original derivation of eq. (31) in §3.1 did not utilize the TDF.
@@
To express constraints from SHS tests under stationary contact, we introduced the NSC in §3.3. However, the definition of the NSC (eq. 46) assumes the TDF; hence, we here need to express the stationary contact in other ways. We do this by replacing ‘NSC’ with the literal definition of stationary contact, ‘$\delta\to 0$’. Repeating §3.3.2 with ‘NSC’ replaced by ‘$\delta\to 0$’, we have
@@
Next, we need to specify the regimes where each constraint applies. Again, we cannot use expressions that necessitate the TDF. As discussed in §4.2, VS tests require (i) $f=f_{\rm VS}$ for $\Omega\gtrsim V_{\rm after}/V_{\rm before}$ probed by existing VS tests. Note that this requirement (i) is not conditioned on $V_{\rm after}$ itself because the VS canon has been confirmed over a wide range of absolute slip rates. Because of (i), the canonical behavior of SHS tests ($f=1$, eq. 123) is allowed only in the $V$-$\theta$ regime sufficiently below the steady state that excludes the $\Omega$ range in the strengthening phase following the negative velocity step. Meanwhile, the SHS canon requires (ii) $f=f_{\rm SHS|\delta\to 0}$ for stationary contact. This time, as mentioned earlier, we cannot express stationary contact as the NSC (eq. 46), which is only definable under the TDF. Instead, we here express stationary contact by invoking constant-$\tau$ SHS tests. If we, upon the start of the hold period, drop the shear stress by much more than $A$ from the level [$\tau_{\rm SS}(V_{\rm prior})$] just before the start of hold, stationary contact is realized and kept for a long time throughout the hold period (§A.1). When we consider a large-stress-drop limit of $e^{-\Delta\tau/A}\to 0$, $\Omega\to 0$ holds for $0<t_{\rm h}<\infty$. Therefore, from (ii), we can state $f=f_{\rm SHS|\delta\to 0}$ when $\Omega\to 0$. Thus, from (i) and (ii), we have
@@
The same discussion as that for eq. (86) (+ eq. 72) in §4.2.2 guarantees evolution laws conforming eq. (124) automatically satisfies the SS constraint (eq. 31). Hence, like eq. (86) (+ eq. 72) under the TDF, but without the TDF assumption this time, eq. (124) is a necessary and sufficient condition to reproduce the three canons simultaneously, except for the range not experimentally probed so far.
@@
Below is a cautious remark regarding the difference between the low-velocity limit $V\to 0$ and the far-below-the-steady-state limit $\Omega\to 0$. Because $V\to 0$ and $\Omega\to 0$ are equivalent unless $\theta\to\infty$, we cannot distinguish $V\to 0$ from $\Omega\to 0$ when we consider constant-$\tau$ SHS tests for finite hold time. Therefore, the SHS condition (ii) used in the above derivation of eq. (124), which we took as a constraint at the low-$\Omega$ limit, might be a condition at the low-$V$ limit in fact. This difference may matter in describing the state evolution in loadpoint-hold SHS tests where both $\theta$ and $V$ change gradually, and the change in $\Omega$ is far slower so that the limits represented by $\Omega\to 0$ and $V\to 0$ can be different. Hence, some ambiguity is left in the above derivation and eq. (124). Such ambiguity did not exist in the derivation of eq. (86) in §4.2.2 because the NSC cannot exist outside the low-$\Omega$ regime (proven at the beginning of §4). This ambiguity is part of the reason why we first showed eq. (86). Yet, experiments so far confirm that the VS canon holds independently of $V$, which singly rejects the emergence of the conditional branch according to the value of $V$, supporting the $\Omega$-dependent switch as expressed in eq. (124). Equation (86) expresses this with a single constant $\beta$. Such $\beta$, as the scale of the aging-slip switch (cf. Appendix F), also exists for eq (124), not limited to the TDF (though $\dot{\theta}=f(V,\theta)$ is still assumed), but $f$ in eq (124) is largely unconstrained over a transitional $\Omega$ range.
@@
To sum up, the aging-slip switch according to $\Omega$ (eq. 124), namely, the existence of a characteristic value of $\Omega$, $\beta$, in eq. (86), is a general conclusion deduced from the three canons as far as $\dot{\theta}=f(V,\theta)$. Excepting Bhattacharya et al. (2015, 2022) discussed later in §5.3, nobody has observed this switch, but there is no evidence denying it so far. Given the generality kept in the derivations, if future experiments reject the existence of $\beta$, it will logically mean that the three canons are not the case somewhere in the regimes experimentally probed so far (again, if we ignore the possibility $\dot{\theta}\neq f(\theta,V)$). Simultaneous reproduction of the three canons is the target of the present work and the historical development of the RSF. As has turned out in the present work, the difficulty of reconciling the aging and slip laws is not such a superficial matter that just nobody has yet thought of a clever formula to reproduce the three canons simultaneously. Thence, we believe the general conclusion of this work, that is, aging-slip switch around $\Omega\sim\beta$, has fundamental importance, warranting future experimental and microphysical studies to verify or falsify it.
@@ 5.2 Macrophysical implications of the $\Omega$-controlled aging-slip switch
Noting $\Omega=\theta/\theta_{\rm SS}$, the $\Omega$-dependent nature of the aging-slip switch means that the switch occurs according to the relative difference of instantaneous $\theta$ from its steady-state value for the instantaneous slip rate. When the strength minimum is incorporated, the switch occurs according to $\tilde{\Omega}/\tilde{\Omega}_{\rm SS}=\tilde{\theta}/\tilde{\theta}_{\rm SS}$, as in $\tilde{\psi}_{\rm c}(\tilde{\Omega}/\tilde{\Omega}_{\rm SS})$ (eq. 116), so that the control factor of the switch becomes the relative difference of $\tilde{\theta}=\theta+\theta_{\rm X}$ from its steady-state value. Then, under Dieterich (1978)’s interpretation of $\theta$ as the average contact time, it is implied the aging-slip switch is controlled by the contact time for $\theta\gg\theta_{\rm X}$ while it becomes almost independent of the contact time for $\theta\ll\theta_{\rm X}$. Such clumsiness, however, disappears in the $\Phi$-notation. From eqs. (15) and (105)–(111), we have
@@
For $\theta,\theta_{\rm SS}\gg\theta_{\rm X}$, from eqs. (6), (7), and (30), we have
@@
Notice that the direct effect does not contribute to the difference of the shear stress from its steady-state value (eq. 7), that is, $\tau-\tau_{\rm SS}=\Phi-\Phi_{\rm SS}$. Hence, for any values of $\theta$, our evolution laws set the aging-slip switch according to $\Phi-\Phi_{\rm SS}$ ($\tau-\tau_{\rm SS}$). The threshold $\Phi-\Phi_{\rm SS}$ value of the aging-slip switch is $B\ln\beta$ (eqs. 86 and 125). Note that eq. (125) also indicates that $\tilde{\Omega}/\tilde{\Omega}_{\rm SS}$ affects strength change only through its logarithm. Then, the crossover regime of the aging-slip switch around $\Omega\sim\beta$ in eq. (86) is narrowed in the $\Phi$ scale so that eq. (86) switches the constraints on the evolution depending on whether the strength $\Phi$ is significantly higher or lower than the threshold value $\Phi_{\rm SS}+B\ln\beta$.
The above-mentioned $\tau-\tau_{\rm SS}=\Phi-\Phi_{\rm SS}$ rewrites eq. (125) into
@@
By using eq. (127), for interfaces with $|A-B|/B\ll 1$ as typically observed in experiments [e.g., $|A-B|/B\simeq 1/30$ reported by Bhattacharya et al. (2015)], we can propose a more physically intelligible interpretation of the $\Omega$-dependent aging-slip switch as below. In the remaining part of this subsection §5.2, we assume $|A-B|/B\ll 1$ unless otherwise noted. Then, the steady-state frictional resistance $\tau_{\rm SS}$ remains close to the reference value $\tau_{*}$ unless hugely extensive range of velocities is considered; we obtain the following relation from the SS canon (eq. 26), using the identity eq. (127):
@@
As mentioned in §4.1, Kato & Tullis (2001) proposed to switch the aging and slip laws according to the slip rate $V$, under the expectation that there must be a critical slip rate $V_{\rm c}$ where the switch occurs. Upon introducing the composite law, Kato & Tullis (2001) said, “To accommodate a friction law to experimental observations of both slide-hold-slide tests and velocity stepping tests, the evolution equation of $\theta$ must be similar to the slip law near the steady state and must approach unity at very small velocities.” [Note: for readability in the present discussion, we have replaced some expressions from the original: state variable $\to\theta$; (3) $\to$ the slip law]. We can then summarize their requirement on the evolution law, here denoted by $f_{\rm KT}$, in a form analogous to eq. (124):
@@
Here, $f_{\rm VS}$ intends, as in eq. (124), the exponential convergence toward the steady state over a constant slip distance (e.g., eq. 43), but does not include the enforcement of steady-state value following the SS canon. We think this is what Kato & Tullis (2001) intended when they mentioned ‘similar to the slip law’ in the above excerpt. Although ‘the slip law’ (eq. 5) enforces both behaviors of VS and SS, it is clear that the latter SS-canon enforcement was not intended by Kato & Tullis (2001), given that their proposed composite law does not follow the SS canon.
Because $V\to 0$ and $\Omega\to 0$ cannot be distinguished for finite $\theta$ in our derivation of eq. (124), we may say that Kato and Tullis’s (2001) idea expressed by eq. (129) is correct from the conceptual viewpoint of the aging-slip switch. The only major failure of the composite law (eq. 74) is the deviation from the SS canon. [Strictly speaking, eq. (129) has one more problem: constraint by $f_{\rm VS}$ needs to be imposed for a broader $\Omega$ range than said in the first line of eq. (129). This problem can be understood if one notices that the aging law, which does not reproduce the VS canon, actually satisfies eq. (129). Nevertheless, we here ignore this shortcoming for this discussion at an upper conceptual level.]